How much lemonade must be added to a gallon of 80% lemonade solution, in order to create a 95% lemonade mix?
Answer by macston(5194)

.
x=amount of 100% lemonade to add
.
0.80(1 gal) + 1.00(x)=0.95(x+1 gal)
0.80 gal + x = 0.95x + 0.95 gal
0.05x = 0.15 gal
x = 3 gal
ANSWER You must add 3 gallons of 100% lemonade.
.
CHECK:
0.80(1 gal) + 1.00(x)=0.95(x+1 gal)
0.80(1 gal)+1.00(3 gal)=0.95(3 gal + 1 gal)
0.80 gal + 3.00 gal=0.95(4 gal)
3.80 gal=3.80 gal